OpenStudy (anonymous):

Apply Euler's method with the step h = 1/2 to find an approximation of y(1), where y is the solution of the initial value problem y'=1+y ; y(0)=0

OpenStudy (anonymous):

i got y = 7/2 approximately

OpenStudy (anonymous):

anyone care to verify for me?

OpenStudy (anonymous):

y[1] = e-1 = 1.7 Is 7/2 = 3.5 considered close?